# VeriStruct: AI-assisted Automated Verification of Data-Structure Modules in Verus

### v2

@@ Abstract
We introduce VeriStruct, a novel framework that extends AI-assisted automated verification from single functions to more complex data structure modules in Verus. VeriStruct employs a planner module to orchestrate the systematic generation of abstractions, type invariants, specifications, and proof code. To address the challenge that LLMs often misunderstand Verus’ annotation syntax and verification-specific semantics, VeriStruct embeds syntax guidance within prompts and includes a repair stage to automatically correct annotation errors.
In an evaluation on eleven Rust data structure modules, VeriStruct succeeds on ten of the eleven, successfully verifying 128 out of 129 functions (99.2%) in total. These results represent an important step toward the goal of automatic AI-assisted formal verification.
@@ 1 Introduction
The prompt and source code of VeriStruct is available on GitHub at https://github.com/ChuyueSun/VeriStruct[artifact]
@@ 2 Preliminaries: Verifying a Data-Structure in Verus
Example 1 (Verified Ring Buffer).
@@
Annotations include preconditions and postconditions for each function. The implicit program requirement is that: if a function is invoked with a program state satisfying the precondition, then it will return (or terminate) in a state satisfying the postcondition. In Verus, one uses the keywords requires (Lines 38–39) and ensures (Lines 40–41, 47–48) for specifying preconditions and postconditions, respectively. The verifier checks that every call site meets the callee’s preconditions and that every function body establishes its postconditions from the assumed preconditions.
@@
Implementation code may include proof blocks embedded in the code. These blocks provide hints to the verifier to help it discharge difficult proof obligations (Lines 50–52 in Fig.˜1).
@@
The view trait in Verus establishes a bridge between the concrete implementation of a data structure and its abstract mathematical representation used in specifications. It requires implementing a view() function, which defines how an instance of a data structure maps to a logical object that the verifier can reason about. While one could theoretically encode a data structure’s state as a Cartesian product of all its fields, defining a custom view offers better abstraction, clarity, and proof efficiency, allowing specifications and proofs to focus on the logical behavior rather than the implementation details. .
@@
A type invariant is a logical formula that all instances of a data structure must satisfy. In Verus, one declares a type invariant using the #[verifier::type_invariant] attribute. We can call use_type_invariant (Line 51) within a proof block to make invariant facts available the current proof context.
Example 3.
@@ 6 Evaluation
We evaluate VeriStruct on a benchmark set comprising eleven Rust data structure modules drawn from the Verus GitHub repository [verus] as well as other open-source repositories. A list of the benchmarks is shown in Table˜1. The benchmarks cover a set of commonly-used data structures, including a ring buffer, a vector implementation with set abstractions and common algorithms, a polymorphic option type, tree-based modules for standalone nodes and a treemap, a bitmap with 64-bit blocks, and concurrent data structures such as a read–write lock and an invariant framework for shared state. Each benchmark contains 5–21 functions to verify, where the number of functions refers to the total number of member methods in the data structure module and the corresponding test functions. In total, our benchmark set includes 129 functions to be verified across all modules.
@@ Results
We summarize the results in Tables˜2 and 5, VeriStruct successfully solves 10 out of 11 benchmarks. For the only unsolved benchmark (Node), VeriStruct still verifies 11 out of 12 functions. In contrast, the baseline solves only 4 out of 11 benchmarks. In terms of the total number of functions verified, VeriStruct verifies 128 out of 129 (99.2%) functions, whereas the baseline verifies only 52. In summary, VeriStruct solves substantially more benchmarks and verifies significantly more functions, demonstrating that our approach improves the capabilities of LLMs to generate correct annotations.
To provide a more fine-grained comparison, we inspect the verification trajectories of VeriStruct across all benchmarks. For each integer $k$, and for each benchmark, we compute the total number of functions fully verified within the first $k$ LLM invocations. We then sum the numbers across all benchmarks for that value of $k$, plot the result, and then increase $k$. Plotting these points creates a visualization that shows the verification progress as a result of increasing LLM calls. We show the result in Fig.˜6, where the x-axis denotes the LLM invocation index $k$, and the y-axis shows the cumulative number of functions verified within the first $k$ invocations across all benchmarks.
@@
We can observe the consistent dominance of VeriStruct over the baseline throughout the verification process. In particular, VeriStruct achieves the largest gain in verified functions during the second invocation (the first is for the planner), highlighting the effectiveness of our planner in selecting a highly effective initial module that verifies a large portion of functions early on. Overall, the experimental results clearly demonstrate that VeriStruct substantially outperforms the baseline, validating the effectiveness of our workflow design.
@@
6.1 Robustness Evaluation
We repeat the eleven-task evaluation five times to quantify robustness. Each run executes the full workflow with a fresh cache, identical sampling parameters, and up to five repair rounds per benchmark, giving 55 independent executions that isolate variance due to stochastic sampling.
Per-Benchmark Outcomes and Aggregate Statistics
Nine benchmarks verify in every run, Bitmap succeeds in three of five runs, and RingBuffer succeeds once. Aggregating those outcomes yields an 89.1% success rate (49/55 executions) with a 3.6% standard deviation across runs. Table˜3 reports that benchmarks consume 21,921 tokens on average with a 2.7% coefficient of variation in the total per-run token budget, while Table˜4 shows the 10,649–38,514 token range for individual benchmarks.
Table 3: Aggregate robustness statistics across the five runs using token counts.
Statistic Success Rate Avg. Tokens / Benchmark Avg. Input Avg. Output
Mean 89.1% 21,921 8,439 13,481
Std. Dev. 3.6% 587 151 512
Coefficient of Variation 2.7% 1.8% 3.8%
Table 4: Token efficiency by benchmark, grouped by reliability tier. “Token/Success” divides the average total tokens by the empirical success probability.
Benchmark Success Rate Avg. Tokens Token/Success
Tier 1: Highly Reliable (100% success)
Transfer 5/5 (100%) 10,649 10,649
Invariants 5/5 (100%) 12,414 12,414
RwLockVstd 5/5 (100%) 12,846 12,846
Option 5/5 (100%) 17,482 17,482
Atomics 5/5 (100%) 19,326 19,326
Node 5/5 (100%) 33,885 33,885
Treemap 5/5 (100%) 38,514 38,514
Vectors 5/5 (100%) 24,394 24,394
SetFromVec 5/5 (100%) 16,368 16,368
Tier 2: Variable Performance (60% success)
Bitmap 3/5 (60%) 23,324 38,873
Tier 3: Challenging (20% success)
RingBuffer 1/5 (20%) 27,250 136,250
Reliability Tiers and Insights
The benchmarks fall into three tiers: nine highly reliable modules that cover core data structures and concurrency primitives, one variable case (Bitmap) that succeeds 60% of the time, and one challenging case (RingBuffer) that rarely converges. Tier 1 problems verify with 10k–39k tokens each, Tier 2 requires roughly 39k tokens per success, and Tier 3 exceeds 136k tokens per success (Table˜4). Because per-run totals vary by under 3% (Table˜3), these token-efficiency gaps mainly reflect inherent difficulty rather than sampling noise, clarifying which cases demand new reasoning techniques beyond repeated attempts.
@@ 8 Conclusion and Future Work
Appendix A Supplementary Material
This appendix provides additional figures and tables that detail verification trajectories and benchmark summaries.
@@ Appendix A Supplementary Material
Syntax Repair Parser and sequence syntax errors Insert missing view() calls or lemmas, balance delimiters, and correct operators/generics.
Type Repair Type mismatches and constructor invariant failures Rewrite expressions, add explicit generics (e.g., None::<T>), and patch missing requires clauses.
Arithmetic Repair Nonlinear reasoning and overflow/underflow Emit by (nonlinear_arith) proofs, assert bounds, and strengthen loop invariants.
Precondition Repair General preconditions, vector bounds, private access Introduce proof blocks, assert length/index constraints, or adjust calls violating permissions.
Postcondition Repair Failing ensures clauses Add exit proofs, tune invariants, or expose ghost accessors for private fields.
Invariant Repair Loop invariant failures pre- and post-loop Assert invariants before loops, propagate conditions, or revise invalid invariants.
Missing Element Repair Absent imports or trait implementations Insert use statements and synthesize required trait methods with contracts.
Mode Repair Mode mismatches and visibility issues Wrap calls in spec/proof blocks, retag functions, or set open/closed attributes.
Old-Self Repair Missing old(self) references Rewrite self. usages inside requires blocks to old(self)..
Assertion Repair Failed assertions and test expectations Add reveals or lemmas (e.g., lemma_seq_subrange_all) and lift tests into postconditions.
Decrease Repair Unsatisfied decreases obligations Prove measure drops, adjust decreases expressions, or update loop variables.
Invariant Removal Redundant invariant invocations Delete unnecessary self.inv() calls when enforced by #[verifier::type_invariant].

### v3

@@ Abstract
We introduce VeriStruct, a novel framework that extends AI-assisted automated verification from single functions to more complex data structure modules in Verus. VeriStruct employs a planner module to orchestrate the systematic generation of abstractions, type invariants, specifications, and proof code. To address the challenge that LLMs often misunderstand Verus’ annotation syntax and verification-specific semantics, VeriStruct embeds syntax guidance within prompts and includes a repair stage to automatically correct annotation errors. In an evaluation on eleven Rust data structure modules, VeriStruct succeeds on ten of the eleven, successfully verifying 128 out of 129 functions (99.2%) in total. These results represent an important step toward the goal of automatic AI-assisted formal verification.
@@ 1 Introduction
The prompt and source code of VeriStruct is available on GitHub [artifact].
@@ 2 Preliminaries: Verifying a Data-Structure in Verus
Example 1(Verified Ring Buffer)
@@
Annotations include preconditions and postconditions for each function. The implicit program requirement is that: if a function is invoked with a program state satisfying the precondition, then it will return (or terminate) in a state satisfying the postcondition. In Verus, one uses the keywords requires (Lines 38–39) and ensures (Lines 40–41, 48–49) for specifying preconditions and postconditions, respectively. The verifier checks that every call site meets the callee’s preconditions and that every function body establishes its postconditions from the assumed preconditions.
@@
Implementation code may include proof blocks embedded in the code. These blocks provide hints to the verifier to help it discharge difficult proof obligations (Lines 51–53 in Fig.˜1).
@@
The view trait in Verus establishes a bridge between the concrete implementation of a data structure and its abstract mathematical representation used in specifications. It requires implementing a view() function, which defines how an instance of a data structure maps to a logical object that the verifier can reason about. While one could theoretically encode a data structure’s state as a Cartesian product of all its fields, defining a custom view offers better abstraction, clarity, and proof efficiency, allowing specifications and proofs to focus on the logical behavior rather than the implementation details.
@@
A type invariant is a logical formula that all instances of a data structure must satisfy. In Verus, one declares a type invariant using the #[verifier::type_invariant] attribute. We can call use_type_invariant (Line 51) within a proof block to make invariant facts available in the current proof context.
Example 3
@@ 6 Evaluation
We evaluate VeriStruct on a benchmark set comprising eleven Rust data structure modules drawn from the Verus GitHub repository [verus] as well as other open-source repositories. A list of the benchmarks is shown in Table˜1. Although the original sources are publicly available, our benchmarks have been substantially modified with more complete specifications, additional methods, and unit tests. We verified that the versions used in our evaluation do not appear verbatim in the o1-2024-12-17 training snapshot; for example, no form of our fully verified RingBuffer appears in the dataset. Furthermore, even if older versions existed in pretraining data, our baseline results demonstrate that LLMs still struggle to synthesize correct Verus specifications without structured guidance—underscoring the necessity of VeriStruct’s workflow. The benchmarks cover a set of commonly-used data structures, including a ring buffer, a vector implementation with set abstractions and common algorithms, a polymorphic option type, tree-based modules for standalone nodes and a treemap, a bitmap with 64-bit blocks, and concurrent data structures such as a read–write lock and an invariant framework for shared state. Each benchmark contains 5–21 functions to verify, where the number of functions refers to the total number of member methods in the data structure module and the corresponding test functions. In total, our benchmark set includes 129 functions to be verified across all modules.
@@
We also compare against Claude Code [claudecode], a state-of-the-art coding agent that can autonomously invoke external tools such as the Verus verifier. We configure Claude Code (using Claude Sonnet 4.5) to iteratively generate annotations and respond to Verus error messages in an agentic loop, analogous to our baseline but with the added capability of tool invocation and autonomous iteration.
@@ Results
We summarize the results in Tables˜2 and 3. VeriStruct successfully solves 10 out of 11 benchmarks. For the only unsolved benchmark (Node), VeriStruct still verifies 11 out of 12 functions. In contrast, the baseline solves only 4 out of 11 benchmarks. In terms of the total number of functions verified, VeriStruct verifies 128 out of 129 (99.2%) functions, whereas the baseline verifies only 52. Claude Code, despite using a more recent model (Claude Sonnet 4.5) and having the ability to autonomously invoke Verus, verifies 102 functions across 8 benchmarks—substantially better than the single-prompt baseline but still short of VeriStruct’s performance. Notably, Claude Code consumes approximately 24k tokens per benchmark on average, while VeriStruct uses only 22k tokens, demonstrating that our structured workflow achieves better results with comparable or lower token cost. In summary, VeriStruct solves substantially more benchmarks and verifies significantly more functions, demonstrating that our approach improves the capabilities of LLMs to generate correct annotations.
To provide a more fine-grained comparison, we inspect the verification trajectories of VeriStruct across all benchmarks. For each integer $k$, and for each benchmark, we compute the total number of functions fully verified within the first $k$ LLM invocations. We then sum the numbers across all benchmarks for that value of $k$, plot the result, and then increase $k$. Plotting these points creates a visualization that shows the verification progress as a result of increasing LLM calls. We show the result in Fig.˜7, where the x-axis denotes the LLM invocation index $k$, and the y-axis shows the cumulative number of functions verified within the first $k$ invocations across all benchmarks.
@@
We can observe the consistent dominance of VeriStruct over the baseline throughout the verification process. In particular, VeriStruct achieves the largest gain in verified functions during the second invocation (the first is for the planner), highlighting the effectiveness of our planner in selecting a highly effective initial module that verifies a large portion of functions early on. Overall, the experimental results clearly demonstrate that VeriStruct substantially outperforms the baseline, validating the effectiveness of our workflow design. Fig.˜6 further shows the per-benchmark verification progression, where each benchmark makes steady progress toward its ground-truth target through iterative refinement.
@@ 8 Conclusion and Future Work
Reinforcement Learning for Verification
Another promising direction is to apply reinforcement learning (RL) to specification inference and proof generation. The Verus verifier provides a natural reward signal: successful verification yields positive feedback, while failed verification attempts with diagnostic messages can guide policy updates. By training models to optimize for verification success, RL could enable more effective exploration of the specification and proof search space, potentially discovering annotation strategies that are difficult to obtain through prompting alone.
Data Availability Statement
The source code of VeriStruct, including all prompts and the benchmark suite used in our evaluation, is publicly available at https://github.com/ChuyueSun/VeriStruct.
Appendix 0.A Supplementary Material
Table˜4 summarizes the specialized repair heuristics used by VeriStruct, the failure classes they target, and representative corrective actions.
@@ Appendix 0.A Supplementary Material
Syntax Parser and sequence syntax errors Insert missing view() calls or lemmas, balance delimiters, correct operators/generics
Type Type mismatches, constructor invariant failures Rewrite expressions, add explicit generics (e.g., None::<T>), patch missing requires clauses
Arithmetic Nonlinear reasoning, overflow/underflow Emit by(nonlinear_arith) proofs, assert bounds, strengthen loop invariants
Precondition Preconditions, vector bounds, private access Introduce proof blocks, assert length/index constraints, adjust permission-violating calls
Postcondition Failing ensures clauses Add exit proofs, tune invariants, expose ghost accessors for private fields
Invariant Loop invariant failures pre-/post-loop Assert invariants before loops, propagate conditions, revise invalid invariants
Missing Element Absent imports or trait implementations Insert use statements, synthesize required trait methods with contracts
Mode Mode mismatches, visibility issues Wrap calls in spec/proof blocks, retag functions, set open/closed attributes
Old-Self Missing old(self) references Rewrite self. in requires blocks to old(self).
Assertion Failed assertions, test expectations Add reveals or lemmas and lift tests into postconditions
Decrease Unsatisfied decreases obligations Prove measure drops, adjust decreases expressions, update loop variables
Inv. Removal Redundant invariant invocations Delete unnecessary self.inv() calls enforced by type invariant attributes